OpenStudy (xx_beast_xx):

What is (-2/3)x(5/4)?

OpenStudy (xx_beast_xx):

Please help me! @Ashleyisakitty

OpenStudy (xx_beast_xx):

This is my last question in math.

OpenStudy (xx_beast_xx):

@shansha1

OpenStudy (xx_beast_xx):

Are you there? @shansha1

OpenStudy (shansha1):

am trying to do it

OpenStudy (xx_beast_xx):

Ok, btw you look pretty!

OpenStudy (shansha1):

thanks

OpenStudy (xx_beast_xx):

:)

OpenStudy (civicsiscool44):

Answers?

OpenStudy (xx_beast_xx):

It is not multiple choise @civicsiscool44

OpenStudy (xx_beast_xx):

"Choice"

OpenStudy (shansha1):

are u suppose it graph it or something

OpenStudy (civicsiscool44):

Do you need it in the decimal form?

OpenStudy (xx_beast_xx):

Nope, just multiplying fractions

OpenStudy (xx_beast_xx):

-2/3 x 5/4

OpenStudy (xx_beast_xx):

No, I need it in fraction form

OpenStudy (xx_beast_xx):

-0.833 is decimal form

OpenStudy (xx_beast_xx):

So what is -0.833 in fraction form?

OpenStudy (civicsiscool44):

-0.83333333 is your answer

OpenStudy (shansha1):

2/3 x 4/5 = 8/15

OpenStudy (civicsiscool44):

\[-\frac{ 5 }{ 6 }\]

OpenStudy (xx_beast_xx):

Thanks! That is what I thought, just checking! Wait, -5/6 or 8/15

OpenStudy (shansha1):

yea

OpenStudy (xx_beast_xx):

Which one?

OpenStudy (xx_beast_xx):

-5/6? IDK

OpenStudy (civicsiscool44):

-5/6

OpenStudy (xx_beast_xx):

Ok, I will see, thx to both of y'all

OpenStudy (xx_beast_xx):

WOW, I got a 60%! I SUCK!

OpenStudy (xx_beast_xx):

I MISSED 4 QUESTIONS

OpenStudy (shansha1):

wow

OpenStudy (xx_beast_xx):

Man, that was the WORST TEST EVER